OpenStudy (anonymous):

does 3x - 7 mean three times a number x fewer than 7

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Fewer BY 7. :D Not THAN!! :D 3x - 7 If x was 5 The it would be 3*5 - 7 = 15-7 = 8 Reduced BY that number!! :D Understood? :)

OpenStudy (anonymous):

its multiple choice so hdifference of three times a number x and seven the difference of seven and three times a number x three times a number x fewer than seven the quotient of three times number x and seven

OpenStudy (anonymous):

the difference

OpenStudy (anonymous):

so is it a the first answer

OpenStudy (anonymous):

not fewer than. less than is a right expression.

OpenStudy (anonymous):

none of them say less than

OpenStudy (anonymous):

I know but is a right he never answered me

zepdrix (zepdrix):

It's a `difference`. In the sentence, which ever term is listed `second` is the one which is being subtracted. Example: the `difference` of `five` and `four times a number x` would be written as, `5` `-` `4x` Does that maybe help you figure it out? :o

OpenStudy (jdoe0001):

x + x + x = 3x thus x + x + x -7 => 3x -7 so, it means, a number added 3 times, or 3 times that number whatever that amount is, then we subtract 7 from it

OpenStudy (anonymous):

7 less than three times a number

OpenStudy (anonymous):

I still think its a because pemdas is involved and multiplication is first then subtraction

OpenStudy (anonymous):

The difference of three times a number x and seven

OpenStudy (anonymous):

no wait nooo c

OpenStudy (anonymous):

three times a number x fewer than seven

zepdrix (zepdrix):

Yah `a` sounds right :) The `difference` of `three times a number x` and `seven` `3x` `-` `7`